A hot air balloon filled with 1.25 x 10^6 L of ideal gas with outdoor temp. At 11 degrees celcius, the air is heated to 121 degrees celcius. WhT is the volume of the air in the balloon after it is heated? Assume none of the gas escapes from the balloon.
Answer #1

V1 = 1.25 x 10^6 L

T1 = 273 + 11= 284 K

T2 = 273 + 121 = 394 K

V2 = ?

from Charle's law

V1/T1 = V2 / T2

1.25 x 10^6 / 294 = V2 / 394

V2 = 1.67 x 10^6

volume of air ballon after heated = 1.67 x 10^6 L